OpenStudy (anonymous):

A contractor employed 210 men to build a house in 60 days. After 12 days , he was joined by 70 more men. In how many days will the remaining work be finished.

OpenStudy (anonymous):

is it 36

OpenStudy (anonymous):

see 210 men were supposed to finish the remaining work in 60-12 =48 days now 210+70 men will finish the work in x days say

OpenStudy (anonymous):

if more person joins the work should be finished early in other words men and time are inversely related hence we have 210/280 =x/48 or x= 210*48/280

Parth (parthkohli):

After 12 days, 1/5 of the work had been completed. 4/5 of the work was left. Now it's pretty obvious that 70 people take \(60 \times 3 = 180\) days because 70 people are 3 times less than 210. So \(70 \times 4 = 280\) people take \(180 \div 4 = 45\) days to complete the same work. But 4/5 of the work is left, so it'd take \(\dfrac{4}{5} \times 45\) days to complete the remaining work.